A and B are 2 sets n(A-B)=32+x , n(B-A)=5x and n(AnB)=x. how to calculate x value

Grade:

3 Answers

jitender lakhanpal
12 years ago

Dear Thamizsh,

 there is a information missing of that of n(AUB)  because without tihs value we can not find the value of x

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender 

thamizh arasi
12 years ago

but in maths book they gave the x value, thats why i asked the solution

Sri Rangan
9 years ago
solution 
given:
n(A-B)=32+x
n(B-A)= 5x
n(AUB) = x
n(A) = n(B)
 
n(A)= n(A-B) + n(AnB) (formula)
n(A) = 32+X+X
n(A) = 32+X,
n(B)= n(B-A)+n(AnB) (formula)
n(B) = 5X+X
n(B) = 6X.
n(A)=n(B) (given)
32+2X= 6X
32+2X-6X=0
32-4X=0
-4X=-32
X=32/4
X=8